Project

General

Profile

Meeting 29-3-2017

Date: 29th Mar 2017
Time: 16:00 UK time
Attendance: Melissa Ed and Adam
Apologies: Paul

Agenda and Minutes

Prev Actions:
Action: Ed, Spare Boards still need temp on RH boards and bias calib on LH boards.

New Actions:

Status of hardware upgrades, spares and essential works: EO CMW
  • cold heads and boards
  • compressor services and spares
  • waveguide swaps
  • helium leak-tighting

Cryo 1 and 2 are now warm and are are being serviced on April 5th. The waveguide swap will happen on the same day if possible.
We have all four absorbers and plan to swap them soon.
Craig has found the problem with the speed at which we burnt through compressors. There was a short on the interlocks for the cryostat which meant that the compressor was permanently on. We can now fix this but it is a bit of a design flaw.

Helical Efficiency Improvements: AD and Round table discussion,
Adam has some tried including some less than perfect tracks but sees very little improvement in the efficiency. He has some misgivings about a least squares fit as it will be to a tan function and hence will have multiple solutions He will discuss with Ken.

Descoped cooling Tracker design

We have discussed having a meeting to discuss the UK works. Melissa will email all.

2. Software: PK, AD, CHunt, CHeidt

  • Performance and Analysis
  • Validation script
  • Geometry,
  • Calibrations, Configurations
  • Monte Carlo
  • Pattern Recognition
  • Kalman
  • Online Displays

Data Taking Summary MAU, EO, PK, AD

http://www.hep.ph.ic.ac.uk/~mgeorge/Files/2016MICEData/

AoB

Next meeting April 12th.

Phone Details: International Toll-Free Dial-in numbers

Access code: 8161978083

Dial-in Numbers:
Country Toll-free Toll
  • Bulgaria 008001194419
  • China 8008703563
  • Italy 800149831 0236009716
  • Japan 006633812547
  • Netherlands 08000203285
  • Russia 81080022671012
  • Switzerland 0800920039 0445807737
  • UK 08082380274 02034333570
  • US 631.621.5253 877.914.2048